1) An athlete is running in 5 races and in each race he has a 70% chance of winning. what is the probability that he will win at least two races?


2) the average number of cars arriving at a particular red light each day is 4. Assuming a poison distribution, calculate the probability that on a given day, less than three cars will arrive at the red light.


1
Expert's answer
2021-06-13T17:35:35-0400

1) Let "X=" the number of winned races: "X\sim Bin(n, p)"

Given "n=5, p=0.7"

"P(X\geq2)=1-P(X=0)-P(X=1)"

"=1-\dbinom{5}{0}(0.7)^0(1-0.7)^{5-0}-\dbinom{5}{1}(0.7)^1(1-0.7)^{5-1}"

"=1-(0.3)^5-5(0.7)(0.3)^4"

"=1-3.8(0.3)^4=0.96922"

The probability that he will win at least two races is 0.96922.


2) Let "X=" the number of cars arriving at a particular red light : "X\sim Po(\lambda)."


"P(X=x)=\dfrac{e^{-\lambda}\cdot\lambda^x}{x!}"

Given "\lambda=4."


"P(X<3)=P(X=0)+P(X=1)+P(X=2)"

"=\dfrac{e^{-4}\cdot4^0}{0!}+\dfrac{e^{-4}\cdot4^1}{1!}+\dfrac{e^{-4}\cdot4^2}{2!}"


"=e^{-4}(1+4+8)\approx0.238103"